# The earliest phases of CNO enrichment in galaxiesthanks: Full Tables 1 and 2 are only available at the CDS via anonymous ftp to cdsarc.u-strasbg.fr(130.79.128.5) or via http://cdsarc.u-strasbg.fr/viz-bin/cat/J/A+A/xxx/Ayyy

### v1

@@ Abstract
Context. The recent detection of nitrogen-enhanced, metal-poor galaxies at high redshift by the James Webb Space Telescope has sparked renewed interest in exploring the chemical evolution of carbon, nitrogen, and oxygen (the CNO elements) at early times, prompting fresh inquiries into their origins.
Aims. The main goal of this paper is to shed light onto the early evolution of the main CNO isotopes in our Galaxy and in young distant systems, such as GN-z11 at $z=10.6$.
@@
Results. By comparing the model predictions with CNO abundance measurements from high-resolution spectroscopy of an homogeneous sample of Galactic halo stars, we first demonstrate that the scatter observed in the metallicity range $-4.5\leq$ [Fe/H] $\leq-1.5$ can be explained by pre-enrichment from Pop III stars that explode as supernovae (SNe) with different initial masses and energies. Then, by exploiting the chemical evolution model, we provide testable predictions for log(C/N), log(N/O), and log(C/O) vs. log(O/H)+12 in MW-like galaxies observed at different cosmic epochs/redshifts. Finally, by calibrating the chemical evolution model to replicate the observed properties of GN-z11, we provide an alternative interpretation of its log(N/O) abundance ratio, demonstrating that its high N content can be reproduced through enrichment from high-mass faint Pop III SNe.
Conclusions. Stochastic chemical enrichment from primordial stars explains both the observed scatter in CNO abundances in MW halo stars and the exceptionally high N/O ratios in some distant galaxies. These findings emphasize the critical role of Pop III stars in shaping early chemical evolution.
@@ 1 Introduction
Comparing chemical abundance measurements in gas and stars with the predictions of galactic chemical evolution (GCE) models allows to place tight constraints on the modes and time scales of the assembly of galactic (sub)structures, as well as on stellar evolution and nucleosynthesis theories. In particular, analyzing the median or mean trends of chemical abundance ratios for elements originating from different nucleosynthetic channels serves as a diagnostic tool for understanding various aspects such as star formation efficiency, gas accretion, outflows, and formation history of galaxies (e.g., Griffith et al., 2019; Spitoni et al., 2019; Weinberg et al., 2019; Matteucci, 2021). At the same time, the dispersion of data points around these mean trends provides further insights. Indeed, the scatter around the mean trends can stem from variations in the relative contributions of different stellar polluters, from the stochastic star formation process itself, or from both. At low metallicities, $\rm[Fe/H]<-2$, a significant increase of the star-to-star scatter is observed for different elements (e.g., Griffith et al., 2023). This evidence has been interpreted as a distinctive indicator of the presence of stochastic enrichment processes (e.g., Argast et al., 2004; Welsh et al., 2021) along with the sign of the chemical enrichment produced by the first stars (e.g., Salvadori et al., 2007; Koutsouridou et al., 2023; Vanni et al., 2023).
@@
On the above theoretical premises, it is clear that there is little hope to observe directly primordial (zero-metal) stars in the local universe: the record holder as the star with the lowest proportion of metals, the Galactic halo star SDSS J102915+172927, is extremely metal-poor ($Z\leq 4.5\times 10^{-5}\,Z_{\odot}$, Caffau et al., 2011a), but no totally free of metals. The chemical footprint of the first stars can be found in the atmospheres of the oldest low-mass stars that have survived to the present day (Cayrel et al., 2004; Beers & Christlieb, 2005; Frebel et al., 2005; de Bennassuti et al., 2014; Salvadori et al., 2019; Koutsouridou et al., 2023; Vanni et al., 2023). Indeed, among ancient very metal-poor stars ($\rm[Fe/H]<-2$), the chemical signatures of Pop III supernovae (SNe) with different explosion energies have been identified (Iwamoto et al., 2005; Ishigaki et al., 2014; Ezzeddine et al., 2019; Placco et al., 2021; Skúladóttir et al., 2021). In the mass range 10–100 M, Pop III SNe can release very different explosion energies, ranging from $\rm E_{SN}\approx 10^{50}$ erg to $\rm E_{SN}\approx 10^{52}$ erg (e.g. Kobayashi et al. 2006; Heger & Woosley 2010), while in between 100 and 140 M they collapse directly to black holes. Conversely, Pop III stars with initial masses within the 140–260 M range are expected to die as pair-instability supernovae (PISNe, Heger & Woosley, 2002), i.e., thermonuclear explosions following the gravitational collapse triggered by dynamical instability of the CO core. The instability is in turn driven by the production and subsequent annihilation of electron-positron pairs in the core, which occurs when low density ($\leq 10^{6}$ g cm-3) and high temperature ($\geq 10^{9}$ K) conditions are met (Rakavy & Shaviv, 1967). The pair-instability collapse leaves no remnant and restores to the interstellar medium (ISM) all the mass of the star, half of which in the form of newly-produced metals. Stars more massive than 260 M, instead, directly collapse to black holes and do not contribute to the chemical enrichment of the surrounding medium (Heger & Woosley, 2002). While including the effects of rotation and magnetic torques may change the mass range in which PISNe are expected to occur (Yoon et al., 2012; Uchida et al., 2019), the crux is that primordial PISNe are expected to imprint a characteristic chemical pattern to the stellar generations that form from their ashes. It has been convincingly demonstrated that long-lived stars born from the gas polluted by the first very massive stars exploding as PISNe should have underabundances of some key elements (N, Cu, Zn) with respect to Fe and [Ca/H] $\simeq-2.5$, [Fe/H] $\simeq-2$ (Karlsson et al., 2008; Salvadori et al., 2019). Due to this metallicity overshoot, PISN descendants should not be searched among the most metal-poor stars. Below [Fe/H]  $\simeq-5$, the almost totality of the stars are expected to form from gas uniquely polluted by faint SNe – the end products of Pop III stars with masses lower than 100 M. Such faint SNe are characterised by large [C/Fe] ratios in the ejecta, a chemical signature that is washed out at higher metallicities because of the increasing contribution of Pop II ‘normal’ core-collapse SNe (CCSNe) as time goes by (e.g., Salvadori et al., 2015; Hartwig et al., 2018).
@@
In this work, we reassess the role of Pop III stars as early Milky Way (MW) pollutants. In particular, we focus on the evolution of the CNO elements, in view of their importance as gas coolants at low metallicities (Bromm & Loeb, 2003; Sharda et al., 2023), alternative H2-gas and star-formation rate tracers (Narayanan & Krumholz, 2014; Combes, 2018; Hunter et al., 2024, and references therein), and constituents of complex organic molecules that possibly precede the formation of prebiotic species (Herbst & van Dishoeck, 2009; Coletta et al., 2020; Fontani et al., 2022; Rocha et al., 2024), not to mention their importance as diagnostics of mixing and nucleosynthetic processes in stars (Charbonnel et al., 1998; Gratton et al., 2000; Lagarde et al., 2024, see also Romano 2022, for a review). Furthermore, we extend our GCE model to high-redshift systems, in which accurate abundances for C, N, and O are being obtained thanks to the Near Infrared Spectrograph (NIRSpec) on board the James Webb Space Telescope (JWST, e.g., Marques-Chaves et al., 2024).
An important, still often overlooked aspect is that of the choice of the proper data set to be used for the comparison with the model outputs. In stellar spectral analyses, for instance, the adoption of different stellar model atmospheres, temperature scales, atomic data, and solar reference abundances introduces systematic offsets that are not all easily identifiable and disposable. Uncritically displaying data from different studies all together most likely leads to an artificial widening of the real underlying dispersion, thus hampering the possibility of a truly meaningful comparison with the model predictions. To avoid that, in this work we derive homogeneous CNO abundances for a sample of bona-fide ‘unmixed’ halo stars (Mucciarelli et al., 2022, see Spite et al. 2005 for a definition of the ‘unmixed’ term), to which we add further data only after a careful evaluation of all systematics. Although it may be superfluous to do so, we reiterate here the importance of comparing GCE model predictions for C and N to abundance measurements of stars that have not yet evolved beyond the lower red giant branch (RGB), thus preserving, in principle, the initial CN abundances in their envelopes. Furthermore, binary stars should be excluded to avoid pollution due to mass transfer from an asymptotic giant branch (AGB) companion star (Bonifacio et al., 2015).
@@
2 Data set
This section introduces the data used to constrain our models for the early Galaxy (Sect. 2.1) and young compact objects detected at high redshift (Sect. 2.2).
@@ 2.1 The Galactic halo
Abundances of CNO elements are homogeneously derived for the sample of lower RGB halo stars presented in Mucciarelli et al. (2022). After a careful evaluation of all possible systematics, we complement this sample with the one of Yong et al. (2013). Abundance estimates for dwarf stars from three-dimensional (3D) radiative transfer calculations including corrections for non local thermodynamic equilibrium (non-LTE) conditions (Amarsi et al., 2019) are added for comparison. We further derive the orbital parameters of the program stars to distinguish those born in situ from the accreted ones (Sect. 2.1.3), and to identify the different progenitor of the latters.
@@ 2.1.1 Stellar Abundances
The main sample used in this work is that presented by Mucciarelli et al. (2022), including 58 RGB stars located before the RGB Bump. Therefore, these stars are not affected by the extra-mixing episode associated to this evolutionary feature (e.g., Cassisi et al., 2002; Karakas & Lattanzio, 2014). For all the targets, Fe, C and Li abundances have been already published (Mucciarelli et al., 2022). For 36 of these targets (covering the metallicity range $\rm-4.5\leq[Fe/H]<-1.0$ dex), the available spectra enabled us to derive N, O and Mg abundances. The spectra used for the analysis were obtained with the Ultraviolet Visual Echelle Spectrograph (UVES, Dekker et al., 2000) at the Very Large Telescope of ESO, details of the data set are in Mucciarelli et al. (2022).
@@
Figure 2: Distribution of the sample stars (circles: stars from Yong et al. 2013, star symbols: stars from Mucciarelli et al. 2022) in the IoM spaces colour-coded according to their $\rm[Fe/H]$ abundances. The APOGEE sample is plotted with filled gray points for reference. The boxes represent the positions occupied by previously identified substructures in the Galactic halo (Dodd et al., 2023).
Figure 3: [C/Fe], [N/Fe], [O/Fe], and [Mg/Fe] vs [Fe/H] abundances for our sample stars. The colour code is same as that for the boxes in Fig.2. Unassociated stars are plotted in gray.
@@
The stellar parameters and abundances for the sample of unmixed halo stars adopted in this paper, which include new homogeneous N and O determinations for the sample of stars presented in Mucciarelli et al. (2022), are listed in Table 1. Throughout this paper, the reference solar abundances are from Grevesse & Sauval (1998) but for O, for which we adopt the photospheric solar abundance determined by Caffau et al. (2011b) from a CO5BOLD 3D model of the solar atmosphere.
@@ 2.1.2 Abundance scatter in CNO elements
Figure 1 reveals a clear trend of increasing mean abundance ratio with decreasing metallicity for all elements (though it should be noticed that all [O/Fe] ratios for [Fe/H] $<-2.5$ are upper limits). At the same time, towards lower metallicities the scatter of measured [X/Fe] abundances, $\Delta\rm[X/Fe]$, increases. For $\rm[Fe/H]>-1.5$, the spread is minimal, with $\Delta\rm[X/Fe]\approx 0$ ($\Delta\rm[X/Fe]\approx 0.6$) for [C/Fe] ([N/Fe]), but only 2 stars fall in this metallicity range. However, in the range $-3.5<\rm[Fe/H]<-2.5$, the scatter increases significantly to $\Delta\rm[X/Fe]\approx 1$ for [C/Fe] and $\Delta\rm[X/Fe]\approx 2$ for [N/Fe]. The observed scatter then apparently decreases for $\rm[Fe/H]<-4$, though it should be emphasized that the sample consists of only four stars within this range. Indeed, adding a sub-sample of data for low-metallicity stars from Yong et al. (2013, circles) leads to larger scatters in both [C/Fe] and [N/Fe] (O measurements are not provided by Yong et al.). Expanding the sample size to include more abundance determinations of unevolved halo stars would allow a better estimate of the dispersion and it is, hence, highly desirable.
@@ 2.1.3 Orbital parameters
As shown in Fig. 3, the abundance ratios of different substructures tend to overlap in the [C/Fe], [N/Fe], [O/Fe], and [Mg/Fe] versus [Fe/H] diagrams. No obvious distinct trends are observed, even though marginal evidence that some Sequoia stars are among the most enhanced in all of the elements exist. Larger samples are clearly required to derive a robust conclusion; still, with the data at hand we can hazard an interpretation of the spread in the elemental abundances as not due to the sum of distinct but tighter patterns, related to individual mergers. The observed scatter at these very low metallicities seems to be due to some intrinsic processes.

### v2

@@ Abstract
Context. The recent detection of super-solar carbon-to-oxygen and nitrogen-to-oxygen abundance ratios in a bunch of metal-poor galaxies at high redshift by the James Webb Space Telescope has sparked renewed interest in exploring the chemical evolution of carbon, nitrogen, and oxygen (the CNO elements) at early times, prompting fresh inquiries into their origins.
Aims. The main goal of this paper is to shed light onto the early evolution of the main CNO isotopes in our Galaxy and in young distant systems, such as GN-z11 at $z=10.6$ and GS-z12 at $z=12.5$.
@@
Results. By comparing the model predictions with CNO abundance measurements from high-resolution spectroscopy of an homogeneous sample of Galactic halo stars, we first demonstrate that the scatter observed in the metallicity range $-4.5\leq$ [Fe/H] $\leq-1.5$ can be explained by pre-enrichment from Pop III stars that explode as supernovae (SNe) with different initial masses and energies. Then, by exploiting the chemical evolution model, we provide testable predictions for log(C/N), log(N/O), and log(C/O) vs. log(O/H)+12 in MW-like galaxies observed at different cosmic epochs/redshifts. Finally, by calibrating the chemical evolution model to replicate the observed properties of GN-z11 and GS-z12, we provide an alternative interpretation of their high N/O and C/O abundance ratio, respectively, demonstrating that an anomalously high N or C content can be reproduced through enrichment from faint Pop III SNe.
Conclusions. Stochastic chemical enrichment from primordial stars explains both the observed scatter in CNO abundances in MW halo stars and the exceptionally high C/O and N/O ratios in some distant galaxies. These findings emphasize the critical role of Pop III stars in shaping early chemical evolution.
@@ 1 Introduction
Comparing chemical abundance measurements in gas and stars with the predictions of galactic chemical evolution (GCE) models allows to place tight constraints on the modes and time scales of the assembly of galactic (sub)structures, as well as on stellar evolution and nucleosynthesis theories. In particular, analyzing the median or mean trends of chemical abundance ratios for elements originating from different nucleosynthetic channels serves as a diagnostic tool for understanding various aspects such as star formation efficiency, gas accretion, outflows, and formation history of galaxies (e.g., Griffith et al., 2019; Mollá et al., 2019; Spitoni et al., 2019, 2023; Weinberg et al., 2019; Hasselquist et al., 2021; Matteucci, 2021). At the same time, the dispersion of data points around these mean trends provides further insights. Indeed, the scatter around the mean trends can stem from variations in the relative contributions of different stellar polluters, from the stochastic star formation process itself, or from both. At low metallicities, $\rm[Fe/H]<-2$, a significant increase of the star-to-star scatter is observed for different elements (e.g., Griffith et al., 2023). This evidence has been interpreted as a distinctive indicator of the presence of stochastic enrichment processes (e.g., Argast et al., 2004; Welsh et al., 2021) along with the sign of the chemical enrichment produced by the first stars (e.g., Salvadori et al., 2007; Koutsouridou et al., 2023; Vanni et al., 2023).
@@
On the above theoretical premises, it is clear that there is little hope to observe directly primordial (zero-metal) stars in the local universe: the record holder as the star with the lowest proportion of metals is the Galactic halo star SDSS J102915+172927 ($Z\leq 4.5\times 10^{-5}\,Z_{\odot}$, Caffau et al., 2011a). Nevertheless, the chemical footprint of the first stars can be found in the atmospheres of the oldest low-mass stars that have survived to the present day (Cayrel et al., 2004; Beers & Christlieb, 2005; Frebel et al., 2005; de Bennassuti et al., 2014; Salvadori et al., 2019; Koutsouridou et al., 2023; Vanni et al., 2023). Indeed, among ancient very metal-poor stars ($\rm[Fe/H]<-2$), the chemical signatures of Pop III supernovae (SNe) with different explosion energies have been identified (Iwamoto et al., 2005; Ishigaki et al., 2014; Ezzeddine et al., 2019; Placco et al., 2021; Skúladóttir et al., 2021). In the initial mass range 10–100 M, Pop III SNe can release very different explosion energies, ranging from $\rm E_{SN}\approx 10^{50}$ erg to $\rm E_{SN}\approx 10^{52}$ erg (e.g. Kobayashi et al. 2006; Heger & Woosley 2010), while in between 100 and 140 M the stars collapse directly to black holes. Conversely, Pop III stars with initial masses within the 140–260 M range are expected to die as pair-instability supernovae (PISNe, Heger & Woosley, 2002), i.e., thermonuclear explosions following the gravitational collapse triggered by dynamical instability of the CO core. The instability is in turn driven by the production and subsequent annihilation of electron-positron pairs in the core, which occurs when low density ($\leq 10^{6}$ g cm-3) and high temperature ($\geq 10^{9}$ K) conditions are met (Rakavy & Shaviv, 1967). The pair-instability collapse leaves no remnant and restores to the interstellar medium (ISM) all the mass of the star, half of which in the form of newly-produced metals. Stars more massive than 260 M, instead, directly collapse to black holes and do not contribute to the chemical enrichment of the surrounding medium (Heger & Woosley, 2002). While including the effects of rotation and magnetic torques may change the mass range in which PISNe are expected to occur (Yoon et al., 2012; Uchida et al., 2019), the crux is that primordial PISNe are expected to imprint a characteristic chemical pattern to the stellar generations that form from their ashes. It has been convincingly demonstrated that long-lived stars born from the gas polluted by the first very massive stars exploding as PISNe should have underabundances of some key elements (N, Cu, Zn) with respect to Fe and [Ca/H] $\simeq-2.5$, [Fe/H] $\simeq-2$ (Karlsson et al., 2008; Salvadori et al., 2019). Due to this metallicity overshoot, PISN descendants should not be searched among the most metal-poor stars. Below [Fe/H]  $\simeq-5$, the almost totality of the stars are expected to form from gas uniquely polluted by faint SNe – the end products of Pop III stars with masses lower than 100 M. Such faint SNe are characterised by large [C/Fe] ratios in the ejecta, a chemical signature that is washed out at higher metallicities because of the increasing contribution of Pop II ‘normal’ core-collapse SNe (CCSNe) as time goes by (e.g., Salvadori et al., 2015; Hartwig et al., 2018).
@@
In this work, we reassess the role of Pop III stars as early Milky Way (MW) pollutants. In particular, we focus on the evolution of the CNO elements, in view of their importance as gas coolants at low metallicities (Bromm & Loeb, 2003; Sharda et al., 2023), alternative H2-gas and star-formation rate tracers (Narayanan & Krumholz, 2014; Combes, 2018; Hunter et al., 2024, and references therein), and constituents of complex organic molecules that possibly precede the formation of prebiotic species (Herbst & van Dishoeck, 2009; Coletta et al., 2020; Fontani et al., 2022; Rocha et al., 2024), not to mention their importance as diagnostics of mixing and nucleosynthetic processes in stars (Charbonnel et al., 1998; Gratton et al., 2000; Lagarde et al., 2024, see also Romano 2022, for a review). Furthermore, we extend our GCE model to high-redshift systems, in which accurate abundances for C, N, and O are being obtained thanks to the Near Infrared Spectrograph (NIRSpec, Jakobsen et al., 2022) on board the James Webb Space Telescope (JWST, see, e.g., Marques-Chaves et al., 2024).
An important, still often overlooked aspect is that of the choice of the proper data set to be used for the comparison with the model outputs. For instance, in stellar spectral analyses the adoption of different stellar model atmospheres, temperature scales, atomic data, and solar reference abundances introduces systematic offsets that are not all easily identifiable and disposable. Uncritically displaying data from different studies all together most likely leads to an artificial widening of the real underlying dispersion, thus hampering the possibility of a truly meaningful comparison with the model predictions. To avoid that, in this work we derive homogeneous CNO abundances for a sample of bona-fide ‘unmixed’ halo stars (Mucciarelli et al., 2022, see Spite et al. 2005 for a definition of the ‘unmixed’ term), to which we add further data only after a careful evaluation of all systematics. Although it may be superfluous to do so, we reiterate here the importance of comparing GCE model predictions for C and N to abundance measurements of stars that have not yet evolved beyond the lower red giant branch (RGB), thus preserving, in principle, the initial CN abundances in their envelopes. Furthermore, binary stars should be excluded to avoid pollution due to mass transfer from an asymptotic giant branch (AGB) companion star (Bonifacio et al., 2015).
@@
2 Data sets
This section introduces the data used to constrain our models for the early Galaxy (Sect. 2.1) and young systems detected at high redshift (Sect. 2.2).
@@ 2.1 The Galactic halo
Abundances of CNO elements are homogeneously derived for the sample of lower RGB halo stars presented in Mucciarelli et al. (2022). After a careful evaluation of all possible systematics, we complement this sample with the one of Yong et al. (2013). Abundance estimates for dwarf stars from three-dimensional (3D) radiative transfer calculations including corrections for non local thermodynamic equilibrium (non-LTE) conditions (Amarsi et al., 2019) are added for comparison. We further derive the orbital parameters of the program stars to distinguish those born in situ from the accreted ones (Sect. 2.1.3) and to identify the different progenitors of the latter.
@@ 2.1.1 Stellar Abundances
The main sample used in this work is that presented by Mucciarelli et al. (2022), including 58 RGB stars located before the RGB Bump. Therefore, these stars are not affected by the extra-mixing episode associated to this evolutionary feature (e.g., Cassisi et al., 2002; Karakas & Lattanzio, 2014). For all the targets, Fe, C and Li abundances have been already published (Mucciarelli et al., 2022). In particular, C abundances have been derived from either the G band or the CH feature at 3143 Å. For 36 of these targets (covering the metallicity range $\rm-4.5\leq[Fe/H]<-1.0$ dex), the available spectra enabled us to derive N, O and Mg abundances. The spectra used for the analysis were obtained with the Ultraviolet Visual Echelle Spectrograph (UVES, Dekker et al., 2000) at the Very Large Telescope of ESO. Details of the data set are in Mucciarelli et al. (2022).
@@
Figure 2: Distribution of the sample stars colour-coded according to their $\rm[Fe/H]$ abundances (circles: stars from Yong et al. 2013, star symbols: stars from Mucciarelli et al. 2022) in the IoM spaces. The APOGEE sample is plotted with filled gray points for reference. The boxes represent the positions occupied by previously identified substructures in the Galactic halo (Dodd et al., 2023).
Figure 3: [C/Fe], [N/Fe], [O/Fe], and [Mg/Fe] vs [Fe/H] abundances for our sample stars. The colour code is the same as that for the boxes in Fig.2. Unassociated stars are plotted in gray.
@@
The stellar parameters and abundances for the sample of unmixed halo stars adopted in this paper, which includes new homogeneous N and O determinations for the sample of stars presented in Mucciarelli et al. (2022), are listed in Table 1. Throughout this paper, the reference solar abundances are from Grevesse & Sauval (1998) but for O, for which we adopt the photospheric solar abundance determined by Caffau et al. (2011b) from a CO5BOLD 3D model of the solar atmosphere.
@@ 2.1.2 Abundance scatter in CNO elements
Figure 1 reveals a clear trend of increasing mean abundance ratio with decreasing metallicity for all elements (though it should be noticed that all [O/Fe] ratios for [Fe/H] $<-2.5$ are upper limits). At the same time, towards lower metallicities the scatter of measured [X/Fe] abundances, $\Delta\rm[X/Fe]$, increases. For $\rm[Fe/H]>-1.5$, the spread is minimal, with $\Delta\rm[X/Fe]\approx 0$ ($\Delta\rm[X/Fe]\approx 0.6$) for [C/Fe] ([N/Fe]), but only 2 stars fall in this metallicity range. However, in the range $-3.5<\rm[Fe/H]<-2.5$, the scatter increases significantly to $\Delta\rm[X/Fe]\approx 1$ for [C/Fe] and $\Delta\rm[X/Fe]\approx 2$ for [N/Fe]. The observed scatter then apparently decreases for $\rm[Fe/H]<-4$, though it should be emphasized that the sample consists of only four stars within this range. Indeed, adding a sub-sample of data for low-metallicity stars from Yong et al. (2013, circles) leads to larger scatters in both [C/Fe] and [N/Fe] (O measurements are not provided by Yong et al.). Expanding the sample size to include more homogeneous abundance determinations of unevolved halo stars would allow a better estimate of the dispersion and it is, hence, highly desirable.
@@ 2.1.3 Orbital parameters
As shown in Fig. 3, the abundance ratios of different substructures tend to overlap in the [C/Fe], [N/Fe], [O/Fe], and [Mg/Fe] versus [Fe/H] diagrams333 Given the paucity of O abundance determinations for [Fe/H] <absent<-< -2.5, we opt to include Mg to highlight the presence of a scatter in the $\alpha$-element abundances. The early evolution of elements other than CNO ones will be discussed in a forthcoming paper.. No obvious distinct trends are observed, even though marginal evidence that some Sequoia stars are among the most enhanced in all of the elements exist. Larger samples are clearly required to derive a robust conclusion; still, with the data at hand we can hazard an interpretation of the spread in the elemental abundances as not due to the sum of distinct but tighter patterns, related to individual mergers. The observed scatter at these very low metallicities seems to be due to some intrinsic processes.